Due to depreciation the marked price of an old car was reduced by $6600 If the original marked price was $32000. Find the
(i) ratio by which the price was reduced and
(ii) rate of cent to the dollar of the reduction

Answer by stanbon(75887)

Due to depreciation the marked price of an old car was reduced by $6600 If the original marked price was $32000. Find the
(i) ratio by which the price was reduced
ratio of reduction = 6000/32000 = 6/32 = 3/16 = 3:16
-------
(ii) rate of cent to the dollar of the reduction
????
-----
Cheers,
Stan H.
------------